OpenStudy (anonymous):

how do i do this? Francois estimates that at least 104 cyclists will be in the Tour de France. if the team leader wants there to be three times as many pack riders as top finishers in the race, what is the least number of pack riders that can be in the stage

OpenStudy (anonymous):

t=number of top finishers p=number of pack riders "at least 104 cyclists..." means t+p >= 104 'three times as many pack riders as top finishers" means 3t=p or t=p/3 to solve, sub p/3 for t in the first equation and solve:\[\frac{p}{3}+p \ge 104\]\[\frac{4}{3}p \ge 104\]\[p \ge 104*\frac{3}{4}\]\[p \ge 78\]

OpenStudy (anonymous):

at least 78 of the riders must be pack riders.